OpenStudy (anonymous):

What is the mass of 2.56 × 10–4 moles of Fe2O3? A. 4.09 × 10–2 g B. 159.6 g C. 6.23 × 105 D. 1.60 × 0–6 g

OpenStudy (chmvijay):

can u tell me molar mas of Fe2O3

OpenStudy (anonymous):

molecular weight of Fe2O3 =56*2 +16*3 =112+48=160

OpenStudy (chmvijay):

1mole of Fe2O3=160 g of Fe2O3 2.56*10^-4 moles of Fe2O3 =2.56*10^-4 *160

OpenStudy (anonymous):

It's A- if anyone else needs to know

OpenStudy (chmvijay):

yup ur right